The Distance Plus partnership has the following capital balances at the beginning of the current year:

Tiger (50% of profits and losses) $ 70,000
Phil (40%) 40,000
Ernie (10%) 55,000

Each of the following questions should be viewed independently.

  1. If Sergio invests $60,000 in cash in the business for a 25 percent interest, what journal entry is recorded? Assume that the bonus method is used.

  2. If Sergio invests $50,000 in cash in the business for a 25 percent interest, what journal entry is recorded? Assume that the bonus method is used.

  3. If Sergio invests $60,000 in cash in the business for a 20 percent interest, what journal entry is recorded? Assume that the goodwill method is used.
Answer #1

(i)

Existing partners' capitals = 70,000 + 40,000 + 55,000

= $165,000

Capital of new partner Sergio = $60,000

Hence, total capital after admission of the new partner = 165,000 + 60,000

= $225,000

New partner Sergio's share = 25%

Hence, new partner's capital = 225,000 x 25%

= $56,250

Sergio invested $60,000 in return for a capital allocation of $56,250. Hence, difference of $3,750 (60,000 - 56,250) will be treated as bonus.

Bonus will be allocated to the existing partners in their profit sharing ratio

Hence, Tiger's bonus share = 3,750 x 50%

= $1,875

Hence, Phil's bonus share = 3,750 x 40%

= $1,500

Hence, Ernie's bonus share = 3,750 x 10%

= $375

Journal

Date Account title Debit Credit
Cash 60,000
Capital - Sergio 56,250
Capital - Tiger 1,875
Capital - Phil 1,500
Capital - Ernie 375

(ii)

Existing partners' capitals = 70,000 + 40,000 + 55,000

= $165,000

Capital of new partner Sergio = $50,000

Hence, total capital after admission of the new partner = 165,000 + 50,000

= $215,000

New partner Sergio's share = 25%

Hence, new partner's capital = 215,000 x 25%

= $53,750

Sergio invested $50,000 in return for a capital allocation of $53,750. Hence, difference of - $3,750 (50,000 - 53,750) will be treated as bonus.

Bonus will be allocated to the existing partners in their profit sharing ratio

Hence, Tiger's bonus share = - 3,750 x 50%

= - $1,875

Hence, Phil's bonus share = - 3,750 x 40%

= - $1,500

Hence, Ernie's bonus share = - 3,750 x 10%

= - $375

Journal

Date Account title Debit Credit
Cash 50,000
Capital - Tiger 1,875
Capital - Phil 1,500
Capital - Ernie 375   
Capital - Sergio 53,750

(iii)

Amount invested by Sergio = $60,000

Sergio's share = 20%

Implied partnership valuation = 60,000 x 100/20

= $300,000

Existing partners' capitals = 70,000 + 40,000 + 55,000

= $165,000

Capital of new partner Sergio = $60,000

Hence, total capital after admission of the new partner = 165,000 + 60,000

= $225,000

Hence, goodwill = 300,000 - 225,000

= $75,000

Goodwill of $75,000 will be distributed among existing partners' in their profit sharing ratio.

Hence, Tiger's goodwill share = 75,000 x 50%

= $37,500

Hence, Phil's goodwill share = 75,000 x 40%

= $30,000

Hence, Ernie's goodwill share = 75,000 x 10%

= $7,500

Journal

Date Account title Debit Credit
Cash 60,000
Goodwill 75,000
Capital - Sergio 60,000
Capital - Tiger 37,500
Capital - Phil 30,000
Capital - Ernie 7,500

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubts. Thanks